### v1

@@ 3.4.3 GTCTV-DPC
4:  Let $\mathcal{X}=\mathcal{Z}_{t}$, $\mathcal{G}_{d}^{0}=\mathcal{G}_{d}^{t_{1}+1}$, $\mathcal{B}_{d}^{0}=\mathcal{B}_{d}^{t_{1}+1}$; // $\mathcal{X}_{t+1}^{\operatorname{B}}=\operatorname{J}_{\tau\operatorname{B}}\left(\mathcal{Z}_{t}\right)$;
5:  for $t_{1}=0:N_{\text{in}}-1$ do
6:   Update $\mathcal{M}^{t_{1}+1}$ by Eq. Eq. 17;
7:   Update $\mathcal{G}_{d}^{t_{1}+1}$ by Eq. Eq. 19 for each $d\in\Gamma$;
8:   $\mathcal{B}_{d}^{t_{1}+1}=\mathcal{B}_{d}^{t_{1}}+\rho_{t_{1}}\left(\nabla_{d}\mathcal{M}^{t_{1}+1}-\mathcal{G}_{d}^{t_{1}+1}\right)$, for each $d\in\Gamma$;
9:   Let $\rho_{t_{1}+1}=\min\{\nu\rho_{t_{1}},10^{10}\}$; // For accelerating convergence.
10:   $\epsilon_{1}=\|\mathcal{M}^{t_{1}+1}-\mathcal{M}^{t_{1}}\|_{F}^{2}\,/\,\|\mathcal{M}^{t_{1}}\|_{F}^{2}$;
11:   if $\epsilon_{1}<\varepsilon$ then
12:    break.
13:   end if
14:  end for
15:  $\mathcal{X}_{t+1}^{\operatorname{B}}=\mathcal{M}^{t_{1}+1}$;
16:  $\mathcal{X}_{t+1}^{\operatorname{C}}=\mathcal{X}_{t+1}^{\operatorname{B}}-\operatorname{D}_{\sigma_{t}}\left(\mathcal{X}_{t+1}^{\operatorname{B}}\right)$; // $\mathcal{X}_{t+1}^{\operatorname{A}}=\operatorname{J}_{\tau\operatorname{A}}\left(2\mathcal{X}_{t+1}^{\operatorname{B}}-z_{t}-\tau\operatorname{C}\left(\mathcal{X}_{t+1}^{\operatorname{B}}\right)\right)$;
17:  $\mathcal{X}_{t+1}^{\operatorname{A}}=\mathcal{P}_{\Omega^{\bot}}\left(2\mathcal{X}_{t+1}^{\operatorname{B}}-\mathcal{Z}_{t}-\tau\alpha\mathcal{X}_{t+1}^{\operatorname{C}}\right)+\mathcal{P}_{\Omega}\left(\mathcal{Y}\right)$;
18:  $\mathcal{Z}_{t+1}=\mathcal{Z}_{t}+\lambda_{t}\left(\mathcal{X}_{t+1}^{\operatorname{A}}-\mathcal{X}_{t+1}^{\operatorname{B}}\right)$; // $\mathcal{Z}_{t+1}=(1-\lambda_{t})\mathcal{Z}_{t}+\lambda_{t}\operatorname{T}\mathcal{Z}_{t}$;
19:  Let $\sigma_{t+1}=\max\{\sigma_{t}\,/\,\nu,10^{-3}\}$; // For accelerating convergence.
20:  $\epsilon=\|\mathcal{X}_{t+1}^{\operatorname{A}}-\mathcal{X}_{t}^{\operatorname{A}}\|_{F}^{2}\,/\,\|\mathcal{X}_{t}^{\operatorname{A}}\|_{F}^{2}$;
21:  if $\epsilon<\varepsilon$ then
22:   break.
23:  end if
24:end for
25: Return $\mathcal{X}_{t+1}^{\operatorname{A}}$.
4 Convergence analysis
@@
In this section, we present a rigorous convergence analysis for GTCTV-DPC (Algorithm 2), formalized in Corollary 4.4. Starting from the general MIP on $V$ in Eq. Eq. 7, we prove that the operator $\operatorname{T}$ in Eq. Eq. 8 is SPC for the extended range $\tau\in(0,4\beta)$, and we provide different admissible choices for the relaxation parameters $\{\lambda_{t}\}_{t\geq 0}$ in Theorem 4.1, extending prior analyses [1, 16].
Theorem 4.1.
@@
Let $\operatorname{A}$ and $\operatorname{B}$ be maximally monotone, and $\operatorname{C}$ be $\beta$-cocoercive in Eq. Eq. 7. Let $\operatorname{T}$ be the operator defined in Eq. Eq. 8 with stepsize $\tau\in(0,4\beta)$. Then, from any initial point $z_{0}\in V$, the iteration $z_{t+1}=(1-\lambda_{t})z_{t}+\lambda_{t}\operatorname{T}z_{t}$ converges weakly to a fixed point of $\operatorname{T}$, where $\{\lambda_{t}\}_{t\geq 0}$ is a real sequence in $[0,1]$ satisfies (i) $\sum_{t=0}^{\infty}\lambda_{t}=\infty$ and (ii) $\sum_{t=0}^{\infty}\lambda_{t}^{2}<\infty$. Moreover, $x_{t}^{\operatorname{B}}$ and $x_{t}^{\operatorname{A}}$ converge weakly to a solution of the MIP defined in Eq. Eq. 7.
For the proof of Theorem 4.1, we rely on Lemma 4.2, following a similar route as Proposition 2.1 in [16]. However, whereas Davis and Yin [16] conclude that $\operatorname{T}$ is $\frac{2\beta}{4\beta-\tau}$-averaged for $\tau\in(0,2\beta)$, we further demonstrate that $\operatorname{T}$ is SPC for a broader range $\tau\in(0,4\beta)$, and provide alternative admissible choices for the relaxation parameters $\{\lambda_{t}\}_{t\geq 0}$.
Lemma 4.2 (Proof in subsection SM1.5 of the supplement).
Let $\operatorname{A}$ and $\operatorname{B}$ be maximally monotone, and $\operatorname{C}$ be $\beta$-cocoercive in Eq. Eq. 7, $\operatorname{T}$ be the operator defined in Eq. Eq. 8 with the stepsize $\tau>0$. Then, for any $x,y\in V$, we have
(20) $\|\operatorname{T}x-\operatorname{T}y\|^{2}\leq\|x-y\|^{2}+\left(\frac{\tau}{2\beta}-1\right)\left\|(\operatorname{Id}-\operatorname{T})x-(\operatorname{Id}-\operatorname{T})y\right\|^{2}.$
Proof 4.3 (Proof of Theorem 4.1).
We proceed in three parts: (1) show that $\operatorname{T}$ is SPC for $\tau\in\left(0,4\beta\right)$, (2) prove weak convergence to a fixed point of $\operatorname{T}$, and (3) verify that the fixed point solves the MIP in Eq. Eq. 7.
Part 1: $\operatorname{T}$ is SPC. Apply Lemma 4.2, for any $x,y\in V$
(21) $\|\operatorname{T}x-\operatorname{T}y\|^{2}\leq\|x-y\|^{2}+\left(\epsilon-1\right)\left\|(\operatorname{Id}-\operatorname{T})x-(\operatorname{Id}-\operatorname{T})y\right\|^{2}.$
Where $\epsilon=\frac{\tau}{2\beta}$. Since $\tau\in(0,4\beta)$, we have $\epsilon\in(0,2)$. When $\epsilon\in(0,1]$, $\epsilon-1\leq 0$, implying that $\operatorname{T}$ is $k$-SPC with any $k\in(0,1)$ by Definition 2.8. When $\epsilon\in(1,2)$, $\epsilon-1\in(0,1)$, so $\operatorname{T}$ is $k^{\prime}$-SPC with parameter $k^{\prime}=\epsilon-1$.
Furthermore, since
$\|(\operatorname{Id}-\operatorname{T})x-(\operatorname{Id}-\operatorname{T})y\|\leq\|x-y\|+\|\operatorname{T}x-\operatorname{T}y\|,$
substituting into Eq. Eq. 21 yields for $\forall x,y\in V$
$\epsilon\|x-y\|^{2}+2(\epsilon-1)\|x-y\|\|\operatorname{T}x-\operatorname{T}y\|+(\epsilon-2)\|\operatorname{T}x-\operatorname{T}y\|^{2}\geq 0,$
which implies
$\|\operatorname{T}x-\operatorname{T}y\|\leq\frac{\epsilon}{2-\epsilon}\|x-y\|.$
Hence, by Definition 2.10, $\operatorname{T}$ is $L$-Lipschitz with $L=\frac{\epsilon}{2-\epsilon}$.
Part 2: Weak convergence to a fixed point. Since $\operatorname{T}$ is $L$-Lipschitz and SPC on $V$, we know that $\operatorname{T}$ is continuous. According to Schauder’s fixed point theorem in [31], there exists at least one fixed point of $\operatorname{T}$.
Moreover, Lemma 10.6 in [12] ensures that if $\operatorname{T}$ is SPC and $\{\lambda_{t}\}_{t\geq 0}$ is a real sequence in $[0,1]$ satisfies (i) $\sum_{t=0}^{\infty}\lambda_{t}=\infty$, (ii) $\sum_{t=0}^{\infty}\lambda_{t}^{2}<\infty$, then from any initial point $z_{0}\in V$, $\{z_{t}\}_{t\geq 0}$ defined by $z_{t+1}=(1-\lambda_{t})z_{t}+\lambda_{t}\operatorname{T}z_{t}$ converges weakly to a fixed point $z^{*}=\operatorname{T}z^{*}$, i.e., $z^{n}\rightharpoonup z^{*}$. Define $x_{t}^{\operatorname{A}}=\operatorname{J}_{\tau\operatorname{A}}(2x_{t}^{\operatorname{B}}-z_{t}-\tau\operatorname{C}(x_{t}^{\operatorname{B}}))$. By Theorem 2.1 in [16], $x_{t}^{\operatorname{B}}=\operatorname{J}_{\tau\operatorname{B}}(z_{t})\rightharpoonup\operatorname{J}_{\tau\operatorname{B}}(z^{*})=x^{*}$, and $x_{t}^{\operatorname{A}}\rightharpoonup\operatorname{J}_{\tau\operatorname{B}}(z^{*})=x^{*}$.
Part 3: Fixed point solves the MIP Eq. Eq. 7. Let $z^{*}=\operatorname{T}z^{*}$ and $x^{*}=\operatorname{J}_{\tau\operatorname{B}}(z^{*})$. From Eq. Eq. 8, $x^{*}=\operatorname{J}_{\tau\operatorname{A}}(2x^{*}-z^{*}-\tau\operatorname{C}(x^{*}))$. Since $\operatorname{J}_{\tau\operatorname{A}}=(\operatorname{Id}+\tau\operatorname{A})^{-1}$,
$(\operatorname{Id}+\tau\operatorname{A})(x^{*})=2x^{*}-z^{*}-\tau\operatorname{C}(x^{*})\iff\tau\operatorname{A}(x^{*})=x^{*}-z^{*}-\tau\operatorname{C}(x^{*}).$
Since $x^{*}=\operatorname{J}_{\tau\operatorname{B}}(z^{*})=(\operatorname{Id}+\tau\operatorname{B})^{-1}(z^{*})$, we have $z^{*}=x^{*}+\tau\operatorname{B}(x^{*})$. Substituting,
$\tau\operatorname{A}(x^{*})+\tau\operatorname{B}(x^{*})+\tau\operatorname{C}(x^{*})=x^{*}-z^{*}-\tau\operatorname{C}(x^{*})+\tau\operatorname{B}(x^{*})+\tau\operatorname{C}(x^{*})=0.$
Thus, $\operatorname{A}(x^{*})+\operatorname{B}(x^{*})+\operatorname{C}(x^{*})=0$, so $x^{*}$ solves Eq. Eq. 7.
@@
Based on Theorem 4.1, we obtain the following Corollary 4.4, which guarantees the global convergence of Algorithm 2.
Corollary 4.4.
@@
Let $\operatorname{D}_{\sigma}$ be $k$-SPC with $k\in(0,1)$, $\tau\in(0,\frac{2-2k}{\alpha})$, and $\{\lambda_{t}\}_{t\geq 0}$ is a real sequence in $[0,1]$ satisfy (i) $\sum_{t=0}^{\infty}\lambda_{t}=\infty$ and (ii) $\sum_{t=0}^{\infty}\lambda_{t}^{2}<\infty$. Let $\operatorname{T}$ be the operator defined in Eq. Eq. 8. Then, from any initial point $\mathcal{Z}_{0}\in\mathbb{V}$, the iteration $\mathcal{Z}_{t+1}=(1-\lambda_{t})\mathcal{Z}_{t}+\lambda_{t}\operatorname{T}\mathcal{Z}_{t}$ generated by Algorithm 2 converges to a fixed point of $\operatorname{T}$. Moreover, $\mathcal{X}_{t}^{\operatorname{B}}$ and $\mathcal{X}_{t}^{\operatorname{A}}$ converge to a solution of the MIP defined in Eq. Eq. 12.
Proof 4.5.
Since $\operatorname{D}_{\sigma}$ is $k$-SPC with $k\in(0,1)$, Lemma 3.4 and Remark 3.5 imply that $\operatorname{C}=\alpha(\operatorname{Id}-\operatorname{D}_{\sigma})$ is $\beta$-cocoercive with $\beta=\frac{1-k}{2\alpha}$. By Theorem 4.1, for any $\tau\in(0,4\beta)=\left(0,\frac{2-2k}{\alpha}\right)$, the sequence $\{\mathcal{Z}_{t}\}_{t\geq 0}$ converges (weakly) to a fixed point of $\operatorname{T}$ from any initial point $\mathcal{Z}_{0}\in\mathbb{V}$. Consequently, the sequences $\mathcal{X}_{t}^{\operatorname{B}}$ and $\mathcal{X}_{t}^{\operatorname{A}}$ converge (weakly) to a solution of the MIP in Eq. Eq. 12. Finally, since $\mathbb{V}$ is finite-dimensional, weak convergence coincides with strong convergence, completing the proof.
Remark 4.6.
The proposed relaxation parameters allow for larger values in early iterations, gradually decreasing as the iterations proceed, as illustrated in Section 5.2. This design contrasts with that in [16], where the relaxation sequence $\{\lambda_{t}\}_{t\geq 0}$ satisfies $\lambda_{t}\in\left(0,\frac{1}{\alpha}\right)$ for all $t\geq 0$, with $\alpha=\frac{2\beta}{4\beta-\tau}$ and $\tau\in(0,2\beta)$, such that $\sum_{t=0}^{\infty}\left(1-\frac{\lambda_{t}}{\alpha}\right)\frac{\lambda_{t}}{\alpha}=+\infty$.
Moreover, based on Lemma 2.9, we can derive an alternative admissible choice for $\{\lambda_{t}\}_{t\geq 0}$ with $\tau\in(2\beta,4\beta)$ with a formulation analogous to [16] and related work such as [1]. From Lemma 4.2, for $\tau\in(2\beta,4\beta)$, $\operatorname{T}$ is $k$-SPC with $k=\frac{\tau-2\beta}{2\beta}$. According to Lemma 2.9, the sequence $\{z_{t}\}_{t\geq 0}$ can be expressed as
$z_{t+1}$ $=(1-\lambda_{t})z_{t}+\lambda_{t}\operatorname{T}z_{t}=(1-\lambda_{t})z_{t}+\lambda_{t}\left(\frac{1}{1-k}\operatorname{N}-\frac{k}{1-k}\operatorname{Id}\right)z_{t}$
$=z_{t}+\frac{\lambda_{t}}{1-k}\left(\operatorname{N}z_{t}-z_{t}\right),$
where $\operatorname{N}$ is nonexpansive. By Theorem 5.15 in [3], when $\lambda_{t}\in[0,1-k]=\left[0,2-\frac{\tau}{2\beta}\right]$ such that $\sum_{t\geq 0}\lambda_{t}\left(1-k-\lambda_{t}\right)=+\infty$, the sequence $\{z_{t}\}_{t\geq 0}$ converages to a fixed point of $\operatorname{T}$.
Noteably, if $\lambda_{t}$ is strictly bounded within (0,1k)(0,1-k) for all $t\geq 0$, i.e., there exists a sufficiently small $\varepsilon>0$ such that $\lambda_{t}\in(\varepsilon,1-k-\varepsilon)$, then $\lambda_{t}\left(1-k-\lambda_{t}\right)\geq\varepsilon\left(1-k-\varepsilon\right)>0$, and thus $\sum_{t\geq 0}\lambda_{t}\left(1-k-\lambda_{t}\right)=+\infty$. This ensures convergence of $\{z_{t}\}$ to a fixed point of $\operatorname{T}$. Nevertheless, adopting this choice typically confines $\lambda$ to relatively small values, which may slow down the convergence rate in practice.
5 Experiments
@@
In this section, we adopt a Bernoulli random sampling scheme across the entire tensor for all experiments, utilizing publicly available datasets. We evaluate the proposed GTCTV-DPC on two data types: multi-dimensional images and spatio-temporal traffic data. A comprehensive comparative study is conducted against state-of-the-art baselines.
5.1 Experimental settings
@@
For MDI completion, we select twelve MSIs from the CAVE [^2] https://cave.cs.columbia.edu/repository/Multispectral dataset [52], and eleven color videos from the YUV [^3] http://trace.eas.asu.edu/yuv dataset. Selected MSIs and color videos are shown in Figures SM1 and SM2 of the supplement. To comprehensively evaluate GTCTV-DPC on MDI completion, we compare against the following methods: low-rank methods FTNN [22] and t-CTV [48]; deep learning methods HIR-Diff [37], LRTFR [32], and DRO-TFF [26]; multi-prior methods DP3LRTC [56], and FBGND [27]. We evaluate MDI completion using Mean Peak Signal-to-Noise Ratio (MPSNR) and Mean Structural Similarity (MSSIM):
@@
$\operatorname{MPSNR}=\frac{1}{n_{4}}\sum_{i_{4}=1}^{n_{4}}\operatorname{PSNR}\left(\mathcal{X}_{\text{out}}^{(:,i_{4})},\mathcal{X}_{\text{ori}}^{(:,i_{4})}\right),\,\operatorname{MSSIM}=\frac{1}{n_{4}}\sum_{i_{4}=1}^{n_{4}}\operatorname{SSIM}\left(\mathcal{X}_{\text{out}}^{(:,i_{4})},\mathcal{X}_{\text{ori}}^{(:,i_{4})}\right),$
@@
$\operatorname{MAPE}=100\times\frac{1}{n}\sum_{i=1}^{n}\frac{|y_{i}-\hat{y_{i}}|}{|y_{i}|},\,\operatorname{RMSE}=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}},$
@@
5.2 Implementation details

### v2

@@ III-D3 GTCTV-DPC
4:  // $\mathcal{X}_{t+1}^{\operatorname{B}}=\operatorname{J}_{\tau\operatorname{B}}\left(\mathcal{Z}_{t}\right)$;
5:  Let $\mathcal{X}=\mathcal{Z}_{t}$, $\mathcal{G}_{d}^{0}=\mathcal{G}_{d}^{t_{1}+1}$, $\mathcal{B}_{d}^{0}=\mathcal{B}_{d}^{t_{1}+1}$;
6:  for $t_{1}=0:N_{\text{in}}-1$ do
7:   Update $\mathcal{M}^{t_{1}+1}$ by Eq. (16);
8:   Update $\mathcal{G}_{d}^{t_{1}+1}$ by Eq. (18), $d\in\Gamma$;
9:   $\mathcal{B}_{d}^{t_{1}+1}=\mathcal{B}_{d}^{t_{1}}+\rho_{t_{1}}\left(\nabla_{d}\mathcal{M}^{t_{1}+1}-\mathcal{G}_{d}^{t_{1}+1}\right)$, $d\in\Gamma$;
10:   // For accelerating convergence.
11:   Let $\rho_{t_{1}+1}=\min\{\nu\rho_{t_{1}},10^{10}\}$;
12:   $\epsilon_{1}=\|\mathcal{M}^{t_{1}+1}-\mathcal{M}^{t_{1}}\|_{F}^{2}\,/\,\|\mathcal{M}^{t_{1}}\|_{F}^{2}$;
13:   if $\epsilon_{1}<\varepsilon$ then
14:     break.
15:   end if
16:  end for
17:  $\mathcal{X}_{t+1}^{\operatorname{B}}=\mathcal{M}^{t_{1}+1}$;
18:  // $\mathcal{X}_{t+1}^{\operatorname{A}}=\operatorname{J}_{\tau\operatorname{A}}\left(2\mathcal{X}_{t+1}^{\operatorname{B}}-z_{t}-\tau\operatorname{C}\left(\mathcal{X}_{t+1}^{\operatorname{B}}\right)\right)$;
19:  $\mathcal{X}_{t+1}^{\operatorname{C}}=\mathcal{X}_{t+1}^{\operatorname{B}}-\operatorname{D}_{\sigma_{t}}\left(\mathcal{X}_{t+1}^{\operatorname{B}}\right)$;
20:  $\mathcal{X}_{t+1}^{\operatorname{A}}=\mathcal{P}_{\Omega^{\bot}}\left(2\mathcal{X}_{t+1}^{\operatorname{B}}-\mathcal{Z}_{t}-\tau\alpha\mathcal{X}_{t+1}^{\operatorname{C}}\right)+\mathcal{P}_{\Omega}\left(\mathcal{Y}\right)$;
21:  // $\mathcal{Z}_{t+1}=(1-\lambda_{t})\mathcal{Z}_{t}+\lambda_{t}\operatorname{T}\mathcal{Z}_{t}$;
22:  $\mathcal{Z}_{t+1}=\mathcal{Z}_{t}+\lambda_{t}\left(\mathcal{X}_{t+1}^{\operatorname{A}}-\mathcal{X}_{t+1}^{\operatorname{B}}\right)$;
23:  // For accelerating convergence.
24:  Let $\sigma_{t+1}=\max\{\sigma_{t}\,/\,\nu,10^{-3}\}$;
25:  $\epsilon=\|\mathcal{X}_{t+1}^{\operatorname{A}}-\mathcal{X}_{t}^{\operatorname{A}}\|_{F}^{2}\,/\,\|\mathcal{X}_{t}^{\operatorname{A}}\|_{F}^{2}$;
26:  if $\epsilon<\varepsilon$ then
27:   break.
28:  end if
29:end for
30:Return $\mathcal{X}_{t+1}^{\operatorname{A}}$.
III-E Convergence analysis
@@
In this section, we present a rigorous convergence analysis for GTCTV-DPC (Algorithm 2), formalized in Corollary 8. Starting from the general MIP on $V$ in Eq. (6), we prove that the operator $\operatorname{T}$ in Eq. (12) is SPC for the extended range $\tau\in(0,4\beta)$, and we provide different admissible choices for the relaxation parameters $\{\lambda_{t}\}_{t\geq 0}$ in Theorem 7, extending prior analyses [67, 63].
Theorem 7 (Proof in Section S.I-E of the supplement).
@@
Let $\operatorname{A}$ and $\operatorname{B}$ be maximally monotone, and $\operatorname{C}$ be $\beta$-cocoercive in Eq. (6). Let $\operatorname{T}$ be the operator defined in Eq. (12) with stepsize $\tau\in(0,4\beta)$. Then, from any initial point $z_{0}\in V$, the iteration $z_{t+1}=(1-\lambda_{t})z_{t}+\lambda_{t}\operatorname{T}z_{t}$ converges weakly to a fixed point of $\operatorname{T}$, where $\{\lambda_{t}\}_{t\geq 0}$ is a real sequence in $[0,1]$ satisfies
$\text{(i) }\sum_{t=0}^{\infty}\lambda_{t}=\infty\text{ and (ii) }\sum_{t=0}^{\infty}\lambda_{t}^{2}<\infty.$
Moreover, $x_{t}^{\operatorname{B}}$ and $x_{t}^{\operatorname{A}}$ converge weakly to a solution of the MIP defined in Eq. (6).
@@
Based on Theorem 7, we obtain the following Corollary 8, which guarantees the global convergence of Algorithm 2.
Corollary 8 (Proof in Section S.I-F of the supplement).
Let $\operatorname{D}_{\sigma}$ be $k$-SPC with $k\in(0,1)$, $\tau\in(0,\frac{2-2k}{\alpha})$, and $\{\lambda_{t}\}_{t\geq 0}$ is a real sequence in $[0,1]$ satisfy
$\text{(i) }\sum_{t=0}^{\infty}\lambda_{t}=\infty\text{ and (ii) }\sum_{t=0}^{\infty}\lambda_{t}^{2}<\infty.$
@@
Let $\operatorname{T}$ be the operator defined in Eq. (12). Then, from any initial point $\mathcal{Z}_{0}\in\mathbb{V}$, the iteration $\mathcal{Z}_{t+1}=(1-\lambda_{t})\mathcal{Z}_{t}+\lambda_{t}\operatorname{T}\mathcal{Z}_{t}$ generated by Algorithm 2 converges to a fixed point of $\operatorname{T}$. Moreover, $\mathcal{X}_{t}^{\operatorname{B}}$ and $\mathcal{X}_{t}^{\operatorname{A}}$ converge to a solution of the MIP defined in Eq. (10).
IV Experiments
@@
In this section, we adopt a Bernoulli random sampling scheme across the entire tensor for all experiments, utilizing publicly available datasets. We evaluate the proposed GTCTV-DPC on two data types: multi-dimensional images and spatio-temporal traffic data. A comprehensive comparative study is conducted against state-of-the-art baselines. The source code is publicly available at the GitHub repository [^1] https://github.com/peterchen96/TensorCompletionMIP.git.
IV-A Experimental settings
@@
For MDI completion, we select twelve MSIs from the CAVE [^2] https://cave.cs.columbia.edu/repository/Multispectral dataset [68], and eleven color videos from the YUV [^3] http://trace.eas.asu.edu/yuv dataset. Selected MSIs and color videos are shown in Figs S.1 and S.2 from the supplement. To comprehensively evaluate GTCTV-DPC on MDI completion, we compare against the following methods: low-rank methods FTNN [21] and t-CTV [25]; deep learning methods HIR-Diff [69], LRTFR [36], and DRO-TFF [37]; multi-prior methods DP3LRTC [38], and FBGND [47]. We evaluate MDI completion using Mean Peak Signal-to-Noise Ratio (MPSNR) and Mean Structural Similarity (MSSIM):
@@
$\operatorname{MPSNR}$ $=\frac{1}{n_{4}}\sum_{i_{4}=1}^{n_{4}}\operatorname{PSNR}\left(\mathcal{X}_{\text{out}}^{(:,i_{4})},\mathcal{X}_{\text{ori}}^{(:,i_{4})}\right),$
$\operatorname{MSSIM}$ $=\frac{1}{n_{4}}\sum_{i_{4}=1}^{n_{4}}\operatorname{SSIM}\left(\mathcal{X}_{\text{out}}^{(:,i_{4})},\mathcal{X}_{\text{ori}}^{(:,i_{4})}\right),$
@@
[…]
$\operatorname{RMSE}$ $=\sqrt{\frac{1}{n}\sum_{i=1}^{n}(y_{i}-\hat{y_{i}})^{2}},$
@@
Figure 3: Completion results for the color video stefan (SR = 0.05), showing the 25th frame of each completed video. The MPSNR is indicated in the upper-left corner of each image.
IV-B Implementation details